# ConGeo: Robust Cross-view Geo-localization across Ground View Variations

### v1

@@ Abstract
Cross-view geo-localization aims at localizing a ground-level query image by matching it to its corresponding geo-referenced aerial view. In real-world scenarios, the task requires accommodating diverse ground images captured by users with varying orientations and reduced field of views (FoVs). However, existing learning pipelines are orientation-specific or FoV-specific, demanding separate model training for different ground view variations. Such models heavily depend on the North-aligned spatial correspondence and predefined FoVs in the training data, compromising their robustness across different settings. To tackle this challenge, we propose ConGeo, a single- and cross-modal Contrastive method for Geo-localization: it enhances robustness and consistency in feature representations to improve a model’s invariance to orientation and its resilience to FoV variations, by enforcing proximity between ground view variations of the same location. As a generic learning objective for cross-view geo-localization, when integrated into state-of-the-art pipelines, ConGeo significantly boosts the performance of three base models on four geo-localization benchmarks for diverse ground view variations and outperforms competing methods that train separate models for each ground view variation.
@@ 1 Introduction
Given an image captured at ground level, cross-view geo-localization aims to determine the location of the image by referring to its corresponding aerial view [1, 2, 3, 4, 5]. As an important auxiliary positioning technique, the task enables noisy GPS correction [6] and offers practical applications in fields such as navigation [7] and autonomous driving [8].
@@
Cross-view geo-localization is often addressed as a retrieval task, where a ground-level image acts as the query and a geo-tagged overhead image as the reference. In real-world applications, it often requires a high generalization ability, to handle diverse ground view image variations, including varying orientations and reduced field of view (FoV). In the past, models were trained and evaluated under idealised settings, where both views are aligned to the North [9, 10, 11] which only covers limited scenarios in the real-world challenges (see North-aligned setting in the left panel of Fig. 1, rows (a) and (b)). Recent works [2, 12] extend existing North-aligned datasets to more challenging settings, cyclically shifting a ground view panorama by a random angle to obtain orientation variations, or further reducing the FoV from ${}^{\circ}$ to ${}^{\circ}$, ${}^{\circ}$, or ${}^{\circ}$ to simulate limited ground view information (see left panel of Fig. 1, Unknown Orientation setting in rows (c) and Limited FoV setting in row (d)). However, they train and evaluate models separately for each setting. This limits the generalization ability of the model in real-world applications, where the FoV and orientation of the query image are often unknown. Moreover, the orientation-specific or FoV-specific training prevents the model from learning a robust feature representation across different settings. Instead, the models are often biased toward training data to capture specific spatial correspondence that is not generalizable across ground view variations. For example, in the left panel of Fig. 1 (rows (a) and (b)), the two aerial views share similar road directions, which can also be found in the corresponding ground views. Such North-aligned spatial correspondence in the training data can serve as a shortcut for the model to improve the specific setting but may sacrifice the orientation invariance and feature consistency towards ground view variations (Section 6).
@@
Figure 1: ConGeo boosts the robustness across the North-aligned setting, and ground view variations: unknown orientation (${}^{\circ}$) and limited field of views (${}^{\circ}$, ${}^{\circ}$, and ${}^{\circ}$). We compare with SEH [13], DSM [2] and SAIG-D [12] and report Top-1 Recall on the CVUSA [9] dataset, one of the geo-localization benchmarks.
@@
To address these challenges, we introduce ConGeo, a Contrastive method for cross-view Geo-localization. ConGeo builds on a contrastive learning approach to enhance consistency in feature representations of ground views under various orientations or FoVs. With ConGeo, we follow the intuition that a robust model should consistently identify the same reference image across ground view variations and that a single model should be robust to all variations expected in real-world applications. To achieve this, we design two contrastive objectives to enforce the proximity between ground view variations and their original representation: a single-modal contrastive objective and a cross-model one. The ground view single-modal contrastive loss works to align ground view variants with the original (North-aligned, full panorama) images, and the aerial view single-modal contrastive loss minimizes the representation disparity between the original aerial image with its augmented counterpart. The cross-modal (ground view and aerial view) contrastive loss further aligns the query image variants with reference aerial images. By disrupting the initial geometric alignment and mitigating potential shortcuts found in the training data, ConGeo compels the model to focus on learning coherent features across different modalities and view variations.
We run extensive experiments on four geo-localization benchmarks, demonstrating that ConGeo outperforms existing methods across ground view variations. Moreover, our analysis highlights two advantages compared to existing methods: 1) Robustness: ConGeo leads to accurate retrieval results across orientation and FoV variations without recurring to models trained specifically for those settings. In the right panel of Fig. 1, we show three state-of-art methods for unknown orientation and limited FoV (SEH [13], DSM [2], SAIG-D [12]): even when trained separately for each specific setting, the resulting performances remain uneven, calling for more robust feature representations. ConGeo consistently outperforms these methods by a large margin when facing ground view orientation and FoV shifts (Table 1), but also for transfer to unseen locations and unseen ground view variations (Table 8). Furthermore, a robustness analysis shows that ConGeo improves the orientation invariance of the base model (Fig. 5), allowing the network to focus on discriminant features for the geo-localization task, and, unlike previous methods, that are consistent across different settings (Fig. 5 and 6). 2) Versatility: As a simple and flexible learning pipeline, ConGeo demonstrates strong versatility, as it can be plugged into different geo-localization pipelines to boost their robustness under different ground view variations (Table 5).
@@ 2.1 Cross-view Geo-localization
Second is geo-localization with variations in orientation and FoV: compared to strictly North-aligned panoramas, ground images with unknown orientations and limited FoV (e.g., captured by a smartphone or car camera) are more readily available, but make the task more challenging. Recent studies [16, 2, 17] gradually draw attention to this more realistic scenario. Specifically, DSM [2] proposed to crop out the query-activated region in the reference image after orientation estimation for accurate matching. Rodrigues et al. [17] customized a pipeline for the retrieval under limited FoVs, where the limited FoV ground and aerial images are cropped from the original image as a form of data augmentation. Despite the steady progress, these methods still face limitations: first, specific FoV information is required for aerial image/feature cropping during both training and testing, while in real-world data this strong prior is unknown, especially during testing. Second, models trained on a specific setting cannot generalize and do not perform satisfactorily on all FoV scenarios. Thus several specialised models are trained for each setting.
@@ 2.2 Contrastive Learning for Geo-localization
Contrastive learning has shown impressive performance for self-supervised and supervised learning in computer vision [18, 19, 20, 21], including image classification [22], object detection [23], and vision-language tasks [24]. In cross-view image geo-localization, the triplet loss [25, 10, 26, 2, 4] has been a standard choice to solve the cross-view image retrieval geo-localization problem in a contrastive way. More recently, InfoNCE [5, 12] has shown to be an efficient objective for the task. However, contrastive learning in these works is limited to aligning feature embeddings between the vanilla ground view and aerial images to match image pairs; while it has shown a powerful ability for building view-invariant and robust representations [19, 21, 18]. In this work, we further leverage contrastive learning by using both single- and cross-modal contrastive objectives, achieving more robust feature representations shared by ground view variants.
@@ 3.1 Overview
Problem statement. Let $\{I_{q}\}$ be a set of query images (North-aligned ground view), and let $\{I_{r}\}$ be a set of reference images (aerial view). For geo-localization under arbitrary orientation and limited FoVs, the query image $I_{q}$ can be cyclically shifted with a random angle and restricted by a specific FoV (e.g., ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$, and ${}^{\circ}$), resulting in its variation $I^{*}_{q}$ that still shows the same location. For each $I^{*}_{q}$, the cross-view geo-localization across ground view variations aims to retrieve its corresponding reference image in $\{I_{r}\}$.
Figure 2: ConGeo’s learning pipeline. For feature representation in the left and right boxes, the North-aligned ground image ($I_{q}$), the transformed ground image ($I^{*}_{q}$), and the aerial view ($I_{r}$) are sent to their respective encoders. Then in the feature space, the single- and cross-modal contrastive learning losses are applied to enforce the proximity of the paired images.
Architecture. The classic architecture for geo-localization consists of a siamese network, based on Convolutional Neural Network (CNN) or Vision Transformer (ViT) for feature encoding [5, 4, 3, 2]. More specifically, a set of query features $Q:=\{q\}$ is obtained by passing query images into the query encoder: $q=\mathrm{Encoder}_{\mathrm{q}}(I_{q})$. Similarly, the set of reference features $R:=\{r\}$ is obtained by $r=\mathrm{Encoder}_{\mathrm{r}}(I_{r})$. To boost the model’s robustness, ConGeo includes ground image variations $*\{I_{q}{{}^{*}}\}$. Such variations are obtained by applying a transformation $T$ to each ground view $I_{q}$. A random orientation angle $\theta$ is first used to horizontally shift the ground image, followed by the application of a FoV angle $\alpha$ to crop the panorama query to a FoV query $I_{q}{{}^{*}}=T_{q}(I_{q}|\theta,\alpha)$.
@@ 3.2 Learning Objectives
Single-modal Contrastive Learning. To enhance the consistency in feature representations between views across alignment and FoV variations, a single-modal contrastive learning objective is designed to produce similar representations between the original images (full panorama, North-aligned) and their transformed counterpart. The single-modal contrastive loss for the ground view is computed as:
@@
A single-modal contrastive loss is also applied to the aerial views. However, unlike the ground view contrastive loss that is applied between ground view variations, the aerial view contrastive loss is designed to enforce the similarity between the same reference image with two random data augmentations, aiming to obtain more robust feature representations [18]. The aerial view single-modal contrastive objective is:
@@
where $r_{i}$ is the feature embedding of the original aerial image, $r_{+}$ denotes the positive one corresponding to $r^{*}$, which is another possible orientation of the same aerial image. $\tau_{r}$ and $\tau_{q}$ are learnable temperature parameters [5].
Cross-modal Contrastive Learning. The cross-modal contrastive learning in ConGeo comprises two alignment objectives: first, the loss of the base method is used [^1] Thanks to its flexible learning objective, ConGeo can be plugged into other existing geo-localization models. In this section, we use Sample4Geo as an example. In the experiments, we use ConGeo with Sample4Geo [5], TransGeo [4], and SAIG-D [12]., which we refer to as the vanilla loss; it is a cross-view alignment loss, where the model learns to match the embeddings of a query image $q$ with the embeddings of its corresponding aerial image $r\in R$. In Sample4Geo [5], for example, the vanilla loss is:
@@
Second, to further reinforce the model’s robustness to ground view variations, we enforce the alignment between transformed ground images and aerial images with a cross-modal contrastive objective:
@@ 4.1 Datasets and Evaluation Metrics
Datasets. Experiments are performed on four cross-view geo-localization datasets (see supplementary material for detailed description). CVUSA [27] contains 35,532 view pairs for training and 8,884 for evaluation. CVACT [10] is split into training, validation, and test sets, where the first two are of the same size as CVUSA while the test set has 92,802 image pairs. VIGOR [11] contains 90,618 satellite images and 105,214 street-view images and can be divided into two train-test splits: same-area and cross-area. The training and test data in the cross-area subset were collected from different cities. University-1652 [28] contains drone, satellite and street view images. The street view images are commonly with unknown orientation and limited FoV, enabling a real-world evaluation under this challenging setting. Detailed dataset descriptions can be found in the supplementary material.
@@ 4.2 Implementation Details
Hyper-parameters and Environment. Loss weights $w_{1}$, $w_{2}$, and $w_{3}$ are empirically set to 0.5, 0.5, and 0.25, corresponding to a higher weight on the single-modal contrastive loss and lower weight on the cross-modal one. For training, the orientation angle $\theta$ is randomly drawn between ${}^{\circ}$ and ${}^{\circ}$ while the FoV angle $\alpha$ is set to ${}^{\circ}$. For evaluation, the orientation angle and FoV angle depend on different settings (see Section 4.3). Experiments comparing different hyper-parameters are provided in the supplementary materials. Besides, following the default setting, for those who used Sample4Geo as the base model, the weights of the query encoder and the reference encoder are shared. For each experiment, the model is trained for 60 epochs with a batch size of 16. We use the AdamW optimizer with an initial learning rate of 0.0001 and a cosine learning rate scheduler. A single NVIDIA GeForce RTX 4090 is used for all the experiments.
@@ 4.3 Experimental settings
- Unknown orientation. When testing, we shift the full-view ground image by a random angle and use it as the query. This setting is also noted as ${}^{\circ}$.
- Limited FoV. We first apply a random cyclical shift to the ground image (same as unknown orientation) and then reduce the FoV to ${}^{\circ}$, ${}^{\circ}$, or ${}^{\circ}$.
@@
Set Methods FoV=$360^{\circ}$ FoV=$180^{\circ}$ FoV=$90^{\circ}$ FoV=$70^{\circ}$
R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1%
CVUSA CVM-NET [25] 16.3 38.9 49.4 88.1 7.4 22.5 32.6 75.4 2.8 10.1 16.7 55.5 2.6 9.3 15.1 21.8
CVFT [29] 23.4 44.4 55.2 86.6 8.1 24.3 34.5 75.2 4.8 14.8 23.2 61.2 3.8 12.4 19.3 55.6
SEH [13] 85.4 93.5 95.8 - 53.7 72.3 79.0 - 16.6 32.2 40.3 - 7.8 18.8 25.6 -
DSM [2] 78.1 89.5 92.9 98.5 48.5 68.5 75.6 93.0 16.2 31.4 39.9 71.1 8.8 19.9 27.3 61.2
L2LTR [30] - - - - 56.7 80.9 87.8 98.0 26.9 50.5 60.4 86.9 14.0 33.1 43.9 77.7
SAIG-D [12] 72.0 90.2 94.0 99.1 52.5 78.1 85.8 97.7 26.7 50.2 59.8 86.6 20.9 41.4 51.2 80.4
Sample4Geo [5] 93.3 97.5 98.0 99.1 84.6 95.9 97.6 99.5 55.1 78.3 85.0 96.6 40.9 65.4 74.1 93.0
ConGeo 96.6 98.9 99.2 99.7 92.3 97.9 98.7 99.7 55.5 75.4 81.5 93.9 49.1 70.8 78.0 93.1
Sample4Geo [5] 16.3 26.1 31.4 51.7 4.1 8.4 11.3 30.4 2.5 6.7 9.8 26.7 1.5 4.6 6.7 20.4
CVACT CVM-NET [25] 13.1 33.9 45.7 81.8 3.9 13.7 21.2 59.2 1.5 5.7 9.6 38.1 1.2 5.0 8.4 34.7
CVFT [29] 26.8 46.9 55.1 81.0 7.1 18.5 26.8 63.9 1.9 6.3 10.5 39.3 1.5 5.1 8.2 34.6
SEH [13] 77.4 88.6 90.9 - 47.7 67.9 74.3 - 13.9 28.4 36.2 - 6.9 16.5 22.3 -
DSM [2] 72.9 85.7 88.9 95.3 49.1 67.8 74.2 89.9 18.1 33.3 40.9 68.7 8.3 20.7 27.1 57.1
Sample4Geo [5] 82.4 90.6 92.3 95.4 58.9 79.8 85.3 95.0 27.9 52.0 62.3 87.0 18.8 40.4 51.0 81.3
ConGeo 83.0 90.6 92.4 96.3 70.3 85.2 88.6 95.1 40.6 62.6 69.8 86.6 24.6 45.3 54.3 80.6
Sample4Geo [5] 12.6 17.5 19.9 31.2 3.4 8.0 10.6 24.5 1.9 5.5 7.9 23.9 1.0 3.2 5.0 16.8
ConGeo 62.6 79.9 84.7 93.9 70.3 85.2 88.6 95.1 34.8 56.9 64.8 83.6 18.5 37.4 46.7 72.5
Table 1: Comparison on the unknown orientation setting and limited FoV setting on CVUSA [27] and CVACT Val [2] datasets. “-” means the score is not provided in the original paper. The best performance is in bold. Results of models trained with FoV-specific images are reported in the white background, while results from a single model without FoV-specialized training are reported with a grey background.

### v2

# https://eceo-epfl.github.io/ConGeo/

@@ Abstract
Equal contribution  ${}^{\text{\textdagger}}$ Corresponding author (xuchangeis@whu.edu.cn)
Cross-view geo-localization aims at localizing a ground-level query image by matching it to its corresponding geo-referenced aerial view. In real-world scenarios, the task requires accommodating diverse ground images captured by users with varying orientations and reduced field of views (FoVs). However, existing learning pipelines are orientation-specific or FoV-specific, demanding separate model training for different ground view variations. Such models heavily depend on the North-aligned spatial correspondence and predefined FoVs in the training data, compromising their robustness across different settings. To tackle this challenge, we propose ConGeo, a single- and cross-view Contrastive method for Geo-localization: it enhances robustness and consistency in feature representations to improve a model’s invariance to orientation and its resilience to FoV variations, by enforcing proximity between ground view variations of the same location. As a generic learning objective for cross-view geo-localization, when integrated into state-of-the-art pipelines, ConGeo significantly boosts the performance of three base models on four geo-localization benchmarks for diverse ground view variations and outperforms competing methods that train separate models for each ground view variation.
@@ 1 Introduction
Given an image captured at ground level, cross-view geo-localization (CVGL) aims to determine the location of the image by referring to its corresponding aerial view [16, 27, 26, 42, 5]. As an important auxiliary positioning technique, the task enables noisy GPS correction [37] and offers practical applications in fields such as navigation [14] and autonomous driving [6].
@@
Cross-view geo-localization is often addressed as a retrieval task, where a ground-level image acts as the query and a geo-tagged overhead image as the reference. In real-world applications, it often requires a high generalization ability, to handle diverse ground view image variations, including varying orientations and reduced field of view (FoV). In the past, models were trained and evaluated under idealised settings, where both views are aligned to the North [33, 17, 44] which only covers limited scenarios in the real-world challenges (see North-aligned setting in the left panel of Fig. 1, rows (a) and (b)). Recent works [27, 45] extend existing North-aligned datasets to more challenging settings, cyclically shifting a ground view panorama by a random angle to obtain orientation variations, or further reducing the FoV from 360 to 70, 90, or 180 to simulate limited ground view information (see left panel of Fig. 1, Unknown Orientation setting in rows (c) and Limited FoV setting in row (d)). However, they train and evaluate models separately for each setting. This limits the generalization ability of the model in real-world applications, where the FoV and orientation of the query image are often unknown. Moreover, the orientation-specific or FoV-specific training prevents the model from representing robust features across different settings. Instead, the models are often biased toward training data to capture specific spatial correspondence that is not generalizable across ground view variations. For example, in the left panel of Fig. 1 (rows (a) and (b)), the two aerial views share similar road directions, which can also be found in the corresponding ground views. Our analysis shows such North-aligned spatial correspondence in the training data can serve as a shortcut for the model to improve the specific setting but may sacrifice the orientation invariance and feature consistency towards ground view variations.
@@
Figure 1: ConGeo boosts the robustness across ground view variations: North-aligned, unknown orientation (FoV=360) and limited field of views (FoV=70, 90, and 180). We compare with SEH [8], DSM [27] and SAIG-D [45] and report Top-1 Recall on the CVUSA [33] dataset, one of the geo-localization benchmarks.
@@
To address this challenge, we introduce ConGeo, a model-agnostic Con-trastive learning pipeline for cross-view Geo-localization. ConGeo follows the intuition that a robust model should capture consistent features for the same location regardless of orientations or FoVs and identify the same reference image across ground view variations. To achieve that, we design two contrastive objectives to enforce the proximity between ground view variations and their original representation: a single-view contrastive objective and a cross-view one. The ground view contrastive loss works to align ground view variants with the original images (North-aligned, full panorama), and the aerial view contrastive loss minimizes the representation disparity between the aerial image and its augmented counterpart. The cross-view (ground view and aerial view) contrastive loss further aligns the query image variants with reference aerial images. By disrupting the initial geometric correspondence and mitigating potential shortcuts found in the training data, ConGeo compels the model to focus on learning coherent features across different modalities and view variations.
We run extensive experiments on four geo-localization benchmarks, demonstrating that ConGeo outperforms existing methods across ground view variations. Moreover, our results highlight three insights: (1) ConGeo empowers a single model to handle various ground view variations, using a model-agnostic learning pipeline. ConGeo consistently outperforms comparison methods under orientation-specific or FoV-specific training by a large margin when facing ground view orientation and FoV shifts (The right panel of Fig. 1, Table 1). Meanwhile, ConGeo demonstrates strong versatility, as it can be plugged into different geo-localization pipelines (Table 5). (2) We demonstrate the advantages of the proposed contrastive learning pipeline over targeted data augmentation when facing ground view variations (Table 3), especially unseen ones (Table 8). (3) Our analysis reveals that the trade-off between the model’s focus on geometric or semantic cues for matching images from different views affects its robustness among variations (Fig. 5 and 6).
@@ 2.1 Cross-view Geo-localization
Second is geo-localization with variations in orientation and FoV: compared to strictly North-aligned panoramas, ground images with unknown orientations and limited FoV (e.g., captured by a smartphone or car camera) are more readily available, but make the task more challenging. Recent studies [43, 27, 25] gradually draw attention to this more realistic scenario. Specifically, DSM [27] proposed to crop out the query-activated region in the reference image after orientation estimation for accurate matching. Rodrigues et al. [25] customized a pipeline for the retrieval under limited FoVs, where the limited FoV images are cropped from the original image as a form of data augmentation for both views. Despite the steady progress, these methods still face limitations: First, specific FoV information is required for aerial image/feature cropping during both training and testing, while in real-world data this strong prior is unknown, especially during testing. Second, models trained on a specific setting cannot generalize and do not perform satisfactorily on all FoV scenarios. Thus several specialised models are trained for each setting.
@@ 2.2 Contrastive Learning for Geo-localization
Contrastive learning has shown impressive performance for self-supervised and supervised learning in computer vision [3, 9, 12, 29], including image classification [31], object detection [35], and vision-language tasks [15]. Existing methods use either cross-modal objectives to align different modalities (e.g., CLIP [24]) or single-modal ones to enhance single-modal representation (e.g., SLIP [21]). In CVGL, the triplet loss [11, 17, 23, 27, 42] has been a standard choice to solve the cross-view image retrieval problem in a contrastive way. More recently, InfoNCE [22, 5, 45] has shown to be an efficient objective for the task. However, contrastive learning in these works is limited to aligning feature embeddings between the vanilla ground view and aerial images to match image pairs; while it has shown a powerful ability for building view-invariant and robust representations [9, 29, 3]. In this work, we further leverage contrastive learning by using both single- and cross-view contrastive objectives, achieving more robust feature representations shared by ground view variants.
@@ 3.1 Overview
Problem Statement. Let $\{I_{q}\}$ be a set of query images (North-aligned ground view), and let $\{I_{r}\}$ be a set of reference images (aerial view). For geo-localization under arbitrary orientation and limited FoVs, the query image $I_{q}$ can be cyclically shifted with a random angle and restricted by a specific FoV (e.g., 70, 90, 180, and 360), resulting in its variation $I^{*}_{q}$ that still shows the same location. For each $I_{q}$ or $I^{*}_{q}$, the cross-view geo-localization aims to retrieve the corresponding reference image in $\{I_{r}\}$.
Figure 2: ConGeo’s learning pipeline. For feature representation in the left and right boxes, the North-aligned ground image ($I_{q}$), the transformed ground image ($I^{*}_{q}$), and the aerial view ($I_{r}$) are sent to their respective encoders. Then in the feature space, the single- and cross-view contrastive learning losses are applied to enforce the proximity of the paired images.
Architecture. The classic architecture for geo-localization consists of a siamese network, based on Convolutional Neural Network (CNN) or Vision Transformer (ViT) for feature encoding [5, 42, 26, 27]. More specifically, a set of query features $Q:=\{q\}$ is obtained by passing query images into the query encoder: $q=\mathrm{Encoder}_{\mathrm{q}}(I_{q})$. Similarly, the set of reference features $R:=\{r\}$ is obtained by $r=\mathrm{Encoder}_{\mathrm{r}}(I_{r})$. To boost the model’s robustness, ConGeo includes ground image variations $\{I_{q}{{}^{*}}\}$. Such variations are obtained by applying a transformation $T$ to each ground view $I_{q}$. A random orientation angle $\theta$ is first used to horizontally shift the ground image, followed by the application of a FoV angle $\alpha$ to crop the panorama query to a FoV query $I_{q}{{}^{*}}=T_{q}(I_{q}|\theta,\alpha)$.
@@ 3.2 Learning Objectives
Single-view Contrastive Learning. To enhance the consistency in feature representations between views across orientations and FoV variations, a single-view contrastive learning objective is designed to produce similar representations between the original images (North-aligned, full panorama) and their transformed counterpart. The loss for the ground view is computed as:
@@
A similar single-view contrastive loss is also applied to the aerial views. However, unlike the ground view contrastive loss that is applied between ground view variations, the aerial view contrastive loss is designed to enforce the similarity between two distorted versions of the same reference image using random data augmentations, aiming to obtain more robust feature representations [3]. The aerial view contrastive objective is:
@@
where $r_{i}$ is the feature embedding of the original aerial image, $r_{+}$ denotes the positive one corresponding to $r^{*}$, which is the feature of another possible augmentation of the same aerial image. $\tau_{r}$ and $\tau_{q}$ are learnable temperature parameters [5].
Cross-view Contrastive Learning. The cross-view contrastive learning in ConGeo comprises two alignment objectives: First, the loss of the base method is used [^1] Thanks to its flexible learning objective, ConGeo can be plugged into other existing geo-localization models. In this section, we use Sample4Geo [5] as an example. In the experiments, we also plug ConGeo into TransGeo [42], and SAIG-D [45]., which we refer to as the vanilla loss; it is a cross-view alignment loss, where the model learns to match the embeddings of a query image $q$ with the embeddings of its corresponding aerial image $r\in R$. In Sample4Geo [5], for example, the vanilla loss is:
@@
Second, to further reinforce the model’s robustness to ground view variations, we enforce the alignment between transformed ground images and aerial images with a cross-view contrastive objective:
@@ 4.1 Datasets and Evaluation Metrics
Datasets. Experiments are performed on four CVGL datasets. CVUSA [39] contains 35,532 view pairs for training and 8,884 for evaluation. CVACT [17] is split into training, validation, and test sets, where the first two are of the same size as CVUSA while the test set has 92,802 image pairs. VIGOR [44] contains 90,618 satellite images and 105,214 street-view images and can be divided into two train-test splits: same-area and cross-area. The training and test data in the cross-area subset were collected from different cities. University-1652 [41] contains drone, satellite and street view images. The street view images are commonly with unknown orientation and limited FoV, enabling a real-world evaluation under this challenging setting. Detailed dataset descriptions can be found in the supplementary material.
@@ 4.2 Implementation Details
Hyper-parameters and Environment. Loss weights $w_{1}$, $w_{2}$, and $w_{3}$ are empirically set to 0.5, 0.5, and 0.25, corresponding to a higher weight on the single-view contrastive loss and lower weight on the cross-view one. For training, the orientation angle $\theta$ is randomly drawn between 0 and 360 while the FoV angle $\alpha$ is set to 180. For evaluation, the orientation angle and FoV angle depend on different settings (see Section 4.3). Experiments comparing different hyper-parameters are provided in the supplementary materials. Besides, following the default setting, for those who used Sample4Geo as the base model, the weights of the query encoder and the reference encoder are shared. For each experiment, the model is trained for 60 epochs with a batch size of 16. We use the AdamW optimizer with an initial learning rate of 0.0001 and a cosine learning rate scheduler. A single NVIDIA GeForce RTX 4090 is used for all the experiments.
@@ 4.3 Experimental settings
- Unknown Orientation. When testing, we shift the full-view ground image by a random angle and use it as the query. This setting is also noted as FoV=360.
- Limited FoV. We first apply a random cyclical shift to the ground image (same as unknown orientation) and then reduce the FoV to 70, 90, or 180.
@@
Set Methods FoV=$360^{\circ}$ FoV=$180^{\circ}$ FoV=$90^{\circ}$ FoV=$70^{\circ}$ Avg.
R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1
CVFT [28] 23.4 44.4 55.2 86.6 8.1 24.3 34.5 75.2 4.8 14.8 23.2 61.2 3.8 12.4 19.3 55.6 10.0
SEH [8] 85.4 93.5 95.8 - 53.7 72.3 79.0 - 16.6 32.2 40.3 - 7.8 18.8 25.6 - 40.9
DSM [27] 78.1 89.5 92.9 98.5 48.5 68.5 75.6 93.0 16.2 31.4 39.9 71.1 8.8 19.9 27.3 61.2 37.9
SAIG-D [45] 72.0 90.2 94.0 99.1 52.5 78.1 85.8 97.7 26.7 50.2 59.8 86.6 20.9 41.4 51.2 80.4 43.0
Sample4Geo [5] 93.3 97.5 98.0 99.1 84.6 95.9 97.6 99.5 55.1 78.3 85.0 96.6 40.9 65.4 74.1 93.0 68.5
ConGeo 96.6 98.9 99.2 99.7 92.3 97.9 98.7 99.7 55.5 75.4 81.5 93.9 49.1 70.8 78.0 93.1 73.4
Sample4Geo[5] 16.3 26.1 31.4 51.7 4.1 8.4 11.3 30.4 2.5 6.7 9.8 26.7 1.5 4.6 6.7 20.4 6.1
Sample4Geo† [5] 93.2 98.2 99.0 99.8 84.6 95.9 97.6 99.5 45.1 64.8 71.3 86.5 28.4 47.1 54.9 75.8 62.8
CVUSA ConGeo 85.2 95.1 96.9 98.9 92.3 97.9 98.7 99.7 55.9 73.2 79.0 90.9 37.1 55.7 62.8 81.4 67.6
CVFT [28] 26.8 46.9 55.1 81.0 7.1 18.5 26.8 63.9 1.9 6.3 10.5 39.3 1.5 5.1 8.2 34.6 9.3
SEH [8] 77.4 88.6 90.9 - 47.7 67.9 74.3 - 13.9 28.4 36.2 - 6.9 16.5 22.3 - 36.5
DSM [27] 72.9 85.7 88.9 95.3 49.1 67.8 74.2 89.9 18.1 33.3 40.9 68.7 8.3 20.7 27.1 57.1 37.1
Sample4Geo [5] 82.4 90.6 92.3 95.4 58.9 79.8 85.3 95.0 27.9 52.0 62.3 87.0 18.8 40.4 51.0 81.3 47.0
ConGeo 83.0 90.6 92.4 96.3 70.3 85.2 88.6 95.1 40.6 62.6 69.8 86.6 24.6 45.3 54.3 80.6 54.6
Sample4Geo[5] 12.6 17.5 19.9 31.2 3.4 8.0 10.6 24.5 1.9 5.5 7.9 23.9 1.0 3.2 5.0 16.8 4.7
Sample4Geo†[5] 77.1 89.8 92.6 97.2 58.9 79.8 85.3 95.0 22.4 44.0 52.7 78.3 12.5 28.5 37.5 67.4 42.7
CVACT ConGeo 62.6 79.9 84.7 93.9 70.3 85.2 88.6 95.1 34.8 56.9 64.8 83.6 18.5 37.4 46.7 72.5 46.6
Table 1: Comparison on the unknown orientation setting and limited FoV setting on CVUSA [39] and CVACT Val [27] datasets. “-” means the score is not provided in the original paper. The best performance is in bold. Results of models trained with FoV-specific images are reported in the white background, while results from a single model without FoV-specialized training are reported with a grey background. † denotes using the same training FoV as ConGeo.